Pamela is 15 years younger than Jiri. The sum of their ages is 25. What is Jiri’s age?
Answer by ikleyn(52781)

.

Jiri + Pamela = 25


Jiri + (Jiri-15) = 25


2*Jiri = 25 + 15 = 40


Jiri = 40/2 = 20.    ANSWER

Solved.